OpenStudy (anonymous):

Use the information on page 135. Include a list of steps you could use to solve the equation for h, and the height of the second hill of the roller coaster.

OpenStudy (flvskidd):

Are you in college?

OpenStudy (anonymous):

no, high school

OpenStudy (flvskidd):

I'm gonna be a junior I cant seem to figure this out.

OpenStudy (flvskidd):

sorry I wasnt much help

OpenStudy (anonymous):

its cool, thanks anyways

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

Step 1: divide by "g" to move it to right side of equation \[195 - h = \frac{v^2}{2 g}\] Step 2: subtract 195 from both sides to move it to right side \[-h = \frac{v^2}{2g} - 195\] Step 3: multiply both sides by "-1" to change the sign of h from neg to pos \[h = -\frac{v^2}{2g} + 195\]